Lars:

What is the volume of the Cylinder? ______𝝅cm3

Lars:

Shadow:

\[V_{c} = \pi r^2 h\] In words, the volume of a cylinder is equal to pi times the radius, squared, times the height

Shadow:

Here you can see that the radius is given at the top, and the height is given on the right. Just input those into the formula and you should be able to solve for the volume. If the question advises for a certain pi value such as 3.14 or asks you to round to a certain place, of course make sure to do that.

Lars:

So do 10cm x 12cm?

Lars:

1 attachment
Shadow:

Sec

Shadow:

I got the same thing

Lars:

@shadow wrote:
I got the same thing
Ohh okay. So 3769.91𝝅cm3 is my final answer?

Shadow:

𝝅cm3 means that you haven't multiplied by pi, since pi is still there

Shadow:

So actually you should just do the radius, exponent, then height part

Lars:

Does that mean 3769.91 x 3.14 then??

Shadow:

No

Shadow:

\[r^2 \times h\] Basically just do r squared times height. Then put pi and cm^3 next to it.

Shadow:

If they want pi in the answer (like next to the numbers) it means they don't want you to multiply it by pi.

Lars:

3769.911183.14cm^3

Shadow:

Just do 10^2 times 12 then put pi next to it

Shadow:

Say a cylinder has a radius of 3 and a height of 4, and they want the answer in:\[\pi cm^3\] I just do: \[3^2 \times 4 = 9 \times 4 = 36\] \[V_{c} = 36 \pi cm^3\]

Lars:

10^2 times 12 would be 1200

Lars:

And you said put pi next to it, so when you say that do you mean 12003.14 or no?

Lars:

@lars wrote:
And you said put pi next to it, so when you say that do you mean 12003.14 or no?
Or if you're asking 1200 x 3.14 than thats = 3768

Shadow:

@lars wrote:
And you said put pi next to it, so when you say that do you mean 12003.14 or no?
No, it would be: \[1200 \pi cm^3\]

Lars:

Oh does that mean 1200Ļ€cm3 is my final answer or am I doing more math ?

Shadow:

Yeah. When they put 𝝅cm3 next to the answer line, they basically said they don't want you to multiply by pi. They want you to just leave it there next to whatever is r^2 times height

Lars:

@shadow wrote:
Yeah. When they put 𝝅cm3 next to the answer line, they basically said they don't want you to multiply by pi. They want you to just leave it there next to whatever is r^2 times height
Ohh I see. Thank you for bearing with me, I know that I'm not the smartest so sorry if I'm struggling to work problems out with

Shadow:

No problem. Thanks for bearing with me. I was kinda in and out being away from my computer.